# Everything everywhere all at once, a probability-based enhanced sampling approach to rare events

### v1

@@ II.5 Calixarene: Host-Guest systems
Moreover, our committor-based simulations allowed us to characterize different reaction channels in a single calculation, which we depict in Fig. 6 and can be identified by visualizing the projection of the Kolmogorov ensemble in the plane defined by $h$ and the water coordination number inside the cavity $V_{2}$ (see SI Fig. S9). As one could expect, the TS configurations are mostly determined by the distance of the ligand from the host molecule, as proved also by the feature relevance analysis (see SI, Fig. S6), which is sensibly shortened with respect to the fully unbound state. However, from $p_{K}$, it can also be seen that there are two contributions to the TSE, differing in the behavior of water and, similar to what is discussed in Sec. II.3, one dominates over the other.
@@
Figure 7: Convergence of $\boldsymbol{\Delta}\mathbf{G}$ Convergence with simulation time of the estimate for the folding energy of chignolin (A) and the binding energy of the G2 ligand to the OAMe octa-acid guest (B). The average estimates from 3 independent simulations for each system are reported as a blue solid line, whereas the uncertainty, computed as the standard deviation over the three replicas, is depicted as a shaded blue region. The reference values are provided as grey dashed lines, and the $0.5\,k_{B}T$ interval around the reference is marked by grey dotted lines. For chignolin, that is the unbiased estimate from Ref. 32, for the calixarene, an estimate obtained using the enhanced sampling setup of Ref. 37.
III Methods
In the following sections, we first provide a general background on the committor and the variational principle it obeys, and we summarize the concepts of the Kolmogorov bias and the related ensemble we have introduced in Ref.1 (Sec. III.1). We then discuss the improved iterative protocol for the self-consistent determination of the committor function (Sec. III.2) and its use to converge free energy calculations (Sec. III.3), highlighting the advancements from our previous work. Kang, Trizio, and Parrinello (2024) In the last section, we schematically summarize the whole iterative procedure to help contextualize all the details (Sec. III.4).
III.1 Background
@@
However, even this approach still presents some serious challenges. In a rare event scenario, the committor is indeed almost constant in the metastable basins, i.e., $q\simeq 0$ in A and $q\simeq 1$ in B, and smoothly interpolates between these two values in the transition region. As a consequence, the gradients that appear in Eq. 5 are strongly localized on the transition region, which, unfortunately, is exactly the region that is difficult to sample in simulations. To bypass this sampling difficulty, in Ref. 1, we introduced the Kolmogorov bias potential, a new enhanced sampling protocol that, by design, focuses the sampling on the TS, which we introduce in the following section.
@@
Such configurations are indeed completely out of reach in standard MD and are difficult to sample even when conventional metadynamics-like enhanced sampling methods are applied. Laio and Parrinello (2002); Invernizzi and Parrinello (2020) Indeed, in that case, the bias potential added to the system aims at filling the metastable basins to promote transitions between them by lowering the relative free energy barriers. However, in the biased energy landscape, the TS region still remains a local maximum whose sampling is thus unfavored compared to the metastable basins (see Fig. 1). In addition, the quality of the TSE sampling with such calculations also depends on the quality of the used CV.
@@
where $\lambda\simeq 1$ is a parameter used to scale the magnitude of the bias. $V_{K}$ depends on the gradients of the committor function, which, as discussed in the previous section, are localized on the TS by construction. As a consequence, in the resulting biased energy landscape $U_{K}(\mathbf{x})=U(\mathbf{x})+V_{K}(\mathbf{x})$, the TS region is stabilized over the rest of the phase space, favoring its sampling (see Fig. 1).
@@
Self-consistent approach for the committor In Ref. 1, we have shown how the variational principle and the sampling of the Kolmogorov ensemble introduced above can be combined in a self-consistent iterative procedure to obtain the committor function with the help of machine learning. In the iterative procedure, we alternate training and sampling cycles until convergence is reached. The first are aimed at improving a neural-network-based variational parametrization of the committor, and the second generate new and progressively better data to learn from under the effect of the TS-oriented Kolmogorov bias. In the following sections, we present an improved version of the procedure that not only allows for a faster convergence through the iterations but also the calculation of the free energy profile of the process.
III.2 Machine learning the committor function
In the method presented in Ref. 1, the committor function is represented as a neural network (NN) $q_{\theta}(\mathbf{x})=q_{\theta}(\mathbf{d}(\mathbf{x}))$ with learnable parameters $\theta$ that takes as inputs a set of physical descriptors $\mathbf{d}(\mathbf{x})$ (see Fig. 1). Such descriptors are functions of the atomic coordinates and can be, for instance, distances or coordination numbers. The optimization of $q_{\theta}^{n}(\mathbf{x})$, at iteration $n$, is based on the minimization of an objective function $L$ that formalizes the variational principle of Eq. 5 and its boundary conditions. In fact, the loss function presents two terms $L=L_{v}+\alpha L_{b}$, relatively scaled by the hyperparameter $\alpha$. The $L_{v}$ term derives from the variational principle and reads
@@ III.2 Machine learning the committor function
In Ref. 1, the $L_{b}$ term was evaluated on a labeled dataset of $N_{A}$ and $N_{B}$ configurations collected at the first iteration by running short, unbiased MD simulations in the two basins. On the other hand, the $L_{v}$ term was computed on the whole dataset available, thus including both the initial unbiased and biased $N^{n}$ configurations from later iterations. This was helpful, as the sampling cycles under the action of the $V_{K}$ bias were focused on the TS region but not optimal, as the unbiased configurations were incorrectly weighted in the dataset. This issue made it necessary to perform more iterations or correct the corresponding weights (i.e., $w_{i}$ when $i\in A,B$) by an estimate of the free energy difference between the two basins, which is not always readily available.
@@
Our new approach gets rid of this slowdown, as at each sampling iteration, we can effectively visit not only the TS region but also both metastable states (see Sec. III.3). In this way, the new configurations cover the whole phase space, and after proper reweighting, the weights $w_{i}$ automatically follow the real Boltzmann distribution. As a consequence, we can thus avoid evaluating the $L_{v}$ term on the unbiased labeled configurations, making the whole procedure more precise and faster. It should also be noted that, in the first iteration, when only unbiased data are available, the contribution from the $L_{v}$ term was almost null also in the original formulation due to the localization of the gradients $|\nabla q(\mathbf{x})|^{2}$ discussed in Sec. III.1. A second consequence of the higher quality of the data thus generated is a more data-efficient protocol. Indeed, as soon as the enhanced sampling becomes effective (i.e., multiple transitions can be observed), to optimize the parameters $\theta$ at each iteration, it is sufficient to use configurations and weights generated in the previous iteration or, at most, from the previous two.
III.3 Extensive sampling along a committor-based CV
@@
As already anticipated in the introduction, we bring two conceptual improvements to the original recipe of Ref. 1. First, we define a physics-informed collective variable (CV) $z(\mathbf{x})$ that is based on the committor and is effective when used with metadynamics-like enhanced sampling approaches. Second, we improve such approaches by combining them with the Kolmogorov bias potential to favor the TS sampling.
@@
Even if the committor has often been thought to be an ideal CV, Berezhkovskii and Szabo (2005); Li and Ma (2014); Ma and Dinner (2005); He, Chipot, and Roux (2022) we already pointed out in the introduction that its straightforward application in an enhanced sampling context is far from optimal due to the high degeneracy of the metastable states along such a variable, as $q(\mathbf{x_{A}})\simeq 0$ and $q(\mathbf{x_{B}})\simeq 1$, and to its sharp behavior in the transition region. This would suggest that a smoother modification of the committor might be an effective CV. A practical way of simply and economically achieving this result comes at zero cost from the NN-based parametrization of the committor proposed in Ref. 1. There, as we have seen in Sec. III.2 and depicted in Fig. 1, the committor is represented as a NN, and a sharp sigmoid activation function $\sigma$ is applied to the output layer of such a network to enforce the correct functional form to the learned function, i.e., $q(\mathbf{x})=\sigma(z(\mathbf{x}))$. It is thus wiser to employ as a CV, not the activated final committor function, but rather the output $z(\mathbf{x})$ of the NN. In fact, this still preserves the physical information encoded in the committor function, but it is also suitable to be used in an enhanced sampling context as it still has finite gradients in the metastable regions, and it is not as sharp in the TS region.
@@
As one can expect, similar numerical issues could also affect the bias potential $V_{K}$ that, in Ref. 1, was computed as a function of the gradients of the committor function $\nabla q(\mathbf{x})$. In fact, this choice is not optimal in practice as, due to the sharp behavior of $q(\mathbf{x})$, such a computation can easily lead to extremely small values below machine precision in the basins where $q(\mathbf{x})$ is flat (see SI, Fig. S1), thus making the bias almost ineffective there. Conveniently, by taking advantage of the change of variable $q\rightarrow z$ and with the help of some elementary algebra (see SI, Sec. S1), one can bypass this issue, writing the bias potential as a function of the smoother and more numerically stable gradients of $z$.
@@
Like any CV, $z$ can be used, in principle, with any CV-based method. However, here, we opt for the On-the-fly Probability Enhanced Sampling (OPES) method Invernizzi and Parrinello (2020), a recent evolution of metadynamics Laio and Parrinello (2002) that arguably represents the state of the art in this field. We already pointed out in Ref. 1 that, in metadynamics-like frameworks, the metastable basins are filled with an external bias potential to reduce the height of free energy barriers and promote transitions, but the TS-region is still found at a local maximum, whose sampling is thus unfavored (see Fig. 1).
@@
As one can expect, the improved sampling of this new approach is also beneficial to the iterative procedure as it greatly simplifies the dataset construction, exploiting the best from the two biasing schemes. From OPES, it inherits a fast-converging bias potential that allows sampling of the relevant part of the phase space. Invernizzi and Parrinello (2022) This facilitates the reweighting, thus guaranteeing a more accurate estimate of the Boltzmann weights and significantly reduces the number of iterations needed for convergence (see Sec. III.2).
On the other hand, from the Kolmogorov bias, we obtain an extensive sampling of the TSE, which, in turn, makes it possible to apply accurately the variational principle (see Sec. III.1).
III.4 Summary of self-consistent iterative procedure
@@
  • Step 1: The NN-based committor $q_{\theta}^{n}(\mathbf{x})$ at iteration $n$ is optimized on the dataset of configurations $\mathbf{x}^{n}$ and weights $w_{i}^{n}$ updated from the previous iterations. To start the procedure, at the first iteration $n=0$, we build such a dataset by running some (short) simulations in the metastable basins and labeling them accordingly. The optimization is based on the variational principle of Eq. 5, minimizing the loss function of Eqs. 8 and 9 (see Sec. III.2).
  • Step 2: We perform biased simulations under the coupled action of the Kolmogorov and OPES bias potentials, $V^{n}_{K}$ and $V^{n}_{\text{OPES}}$. The latter is applied along the committor-related CV $z^{n}(\mathbf{x})$ and drives the sampling across the whole relevant configurational space between the two states, whereas the former guarantees a thorough sampling of the TS region crucial for applying Eq. 5 (see Sec. III.3).
@@
IV Conclusions
Nonetheless, there is still room for improvement. For example, it would be highly desirable to remove the need for physical descriptors, thus providing an even more general approach, for instance, by exploiting more sophisticated models based on graph neural networks that recently have been making their way also in the enhanced sampling field. Zhang et al. (2024); Dietrich et al. (2024); Zou and Tiwary (2024) On the other hand, from the theoretical point of view, improvements could be obtained by eventually extending the ideas behind the committor function to a more general framework that could help tackle more efficiently complex systems characterized by rugged energetic landscapes and multiple intermediates.
Code and Data Availability
The code for the training of the committor is available through the open-source mlcolvar library Bonati et al. (2023) alongside didactic tutorials. Training and simulation data and inputs are available on GitHub [^1] https://github.com/EnricoTrizio/committor_2.0 alongside the PLUMED Tribello et al. (2014); consortium (2019) interface for the application of the bias.

### v4

@@ II.5 Calixarene: Host-Guest systems
Moreover, our committor-based simulations allowed us to characterize different reaction channels in a single calculation, which we depict in Figure 6 and can be identified by visualizing the projection of the Kolmogorov ensemble in the plane defined by $h$ and the water coordination number inside the cavity $V_{2}$ (see Supplementary Figure S9). As one could expect, the TS configurations are mostly determined by the distance of the ligand from the host molecule, as proved also by the feature relevance analysis (see Supplementary Figure S6), which is sensibly shortened with respect to the fully unbound state. However, from $p_{K}$, it can also be seen that there are two contributions to the TSE, differing in the behavior of water and, similar to what is discussed in Section II.3, one dominates over the other.
@@
Figure 7: Convergence of $\boldsymbol{\Delta}\mathbf{G}$ Convergence with simulation time of the estimate for the folding energy of chignolin (A) and the binding energy of the G2 ligand to the OAMe octa-acid guest (B). The average estimates from independent simulations for each system (3 for chignolin, 4 for OAMe-G2) are reported as a blue solid line, whereas the uncertainty, computed as the standard deviation over the three replicas, is depicted as a shaded blue region. The reference values are provided as grey dashed lines, and the $0.5\,k_{B}T$ interval around the reference is marked by grey dotted lines. For chignolin, that is the unbiased estimate from Ref. 33, for the calixarene, an estimate obtained using the enhanced sampling setup of Ref. 38.
III Discussion
Nonetheless, there is still room for improvement. For example, it would be highly desirable to remove the need for physical descriptors, thus providing an even more general approach, for instance, by exploiting more sophisticated models based on graph neural networks that recently have been making their way also in the enhanced sampling field. Zhang et al. (2024); Dietrich et al. (2024); Zou and Tiwary (2024) On the other hand, from the theoretical point of view, improvements could be obtained by eventually extending the ideas behind the committor function to a more general framework that could help tackle more efficiently complex systems characterized by rugged energetic landscapes and multiple intermediates. This should also include the possibility of estimating kinetic rates directly from the committor as prescribed in transition path theory.Weinan and Vanden-Eijnden (2010) In addition, it would also be interesting to investigate the applicability and possible advantages of employing our committor-related CV with other enhanced sampling methods in which directly using the committor was problematic. Rotskoff, Mitchell, and Vanden-Eijnden (2022); Aristoff et al. (2023)
IV Methods
In the following sections, we first provide a general background on the committor and the variational principle it obeys, and we summarize the concepts of the Kolmogorov bias and the related ensemble we have introduced in Ref.3 (Section IV.1). We then discuss the improved iterative protocol for the self-consistent determination of the committor function (Section IV.2) and its use to converge free energy calculations (Section IV.3), highlighting the advancements from our previous work. Kang, Trizio, and Parrinello (2024) In the last section, we schematically summarize the whole iterative procedure to help contextualize all the details (Section IV.4).
IV.1 Background
@@
However, even this approach still presents some serious challenges. In a rare event scenario, the committor is indeed almost constant in the metastable basins, that is, $q\simeq 0$ in A and $q\simeq 1$ in B, and smoothly interpolates between these two values in the transition region. As a consequence, the gradients that appear in Eq. 5 are strongly localized on the transition region, which, unfortunately, is exactly the region that is difficult to sample in simulations. To bypass this sampling difficulty, in Ref. 3, we introduced the Kolmogorov bias potential, a new enhanced sampling protocol that, by design, focuses the sampling on the TS, which we introduce in the following section.
@@
Such configurations are indeed completely out of reach in standard MD and are difficult to sample even when conventional metadynamics-like enhanced sampling methods are applied. Laio and Parrinello (2002); Invernizzi and Parrinello (2020) Indeed, in that case, the bias potential added to the system aims at filling the metastable basins to promote transitions between them by lowering the relative free energy barriers. However, in the biased energy landscape, the TS region still remains a local maximum whose sampling is thus unfavored compared to the metastable basins (see Figure 1). In addition, the quality of the TSE sampling with such calculations also depends on the quality of the used CV.
@@
where $\lambda\simeq 1$ is a parameter used to scale the magnitude of the bias. $V_{K}$ depends on the gradients of the committor function, which, as discussed in the previous section, are localized on the TS by construction. As a consequence, in the resulting biased energy landscape $U_{K}(\mathbf{x})=U(\mathbf{x})+V_{K}(\mathbf{x})$, the TS region is stabilized over the rest of the phase space, favoring its sampling (see Figure 1).
@@
Self-consistent approach for the committor In Ref. 3, we have shown how the variational principle and the sampling of the Kolmogorov ensemble introduced above can be combined in a self-consistent iterative procedure to obtain the committor function with the help of machine learning. In the iterative procedure, we alternate training and sampling cycles until convergence is reached. The first are aimed at improving a neural-network-based variational parametrization of the committor, and the second generate new and progressively better data to learn from under the effect of the TS-oriented Kolmogorov bias. In the following sections, we present an improved version of the procedure that not only allows for a faster convergence through the iterations but also the calculation of the free energy profile of the process.
IV.2 Machine learning the committor function
In the method presented in Ref. 3, the committor function is represented as a neural network (NN) $q_{\theta}(\mathbf{x})=q_{\theta}(\mathbf{d}(\mathbf{x}))$ with learnable parameters $\theta$ that takes as inputs a set of physical descriptors $\mathbf{d}(\mathbf{x})$ (see Figure 1). Such descriptors are functions of the atomic coordinates, for instance, distances or coordination numbers, and as for the case of machine-learning CVs based on the same framework, Bonati et al. (2023); Trizio et al. (2024) they should be chosen so that they can describe the relevant physics of the systems when considered collectively. In addition, in Ref.3, we have shown how the quality of different descriptor sets can be evaluated using the lowest achievable variational bound as a figure of merit.
The optimization of $q_{\theta}^{n}(\mathbf{x})$, at iteration $n$, is based on the minimization of an objective function $L$ that formalizes the variational principle of Eq. 5 and its boundary conditions. In fact, the loss function presents two terms $L=L_{v}+\alpha L_{b}$, relatively scaled by the hyperparameter $\alpha$. The $L_{v}$ term derives from the variational principle and reads
@@ IV.2 Machine learning the committor function
In Ref. 3, the $L_{b}$ term was evaluated on a labeled dataset of $N_{A}$ and $N_{B}$ configurations collected at the first iteration by running short, unbiased MD simulations in the two basins. On the other hand, the $L_{v}$ term was computed on the whole dataset available, thus including both the initial unbiased and biased $N^{n}$ configurations from later iterations. This was helpful, as the sampling cycles under the action of the $V_{K}$ bias were focused on the TS region but not optimal, as the unbiased configurations were incorrectly weighted in the dataset. This issue made it necessary to perform more iterations or correct the corresponding weights (that is, $w_{i}$ when $i\in A,B$) by an estimate of the free energy difference between the two basins, which is not always readily available.
@@
Our new approach gets rid of this slowdown, as at each sampling iteration, we can effectively visit not only the TS region but also both metastable states (see Section IV.3). In this way, the new configurations cover the whole phase space, and after proper reweighting, the weights $w_{i}$ automatically follow the real Boltzmann distribution. As a consequence, we can thus avoid evaluating the $L_{v}$ term on the unbiased labeled configurations, making the whole procedure more precise and faster. It should also be noted that, in the first iteration, when only unbiased data are available, the contribution from the $L_{v}$ term was almost null also in the original formulation due to the localization of the gradients $|\nabla q(\mathbf{x})|^{2}$ discussed in Section IV.1. A second consequence of the higher quality of the data thus generated is a more data-efficient protocol. Indeed, as soon as the enhanced sampling becomes effective (that is, multiple transitions can be observed), to optimize the parameters $\theta$ at each iteration, it is sufficient to use configurations and weights generated in the previous iteration or, at most, from the previous two.
IV.3 Extensive sampling along a committor-based CV
@@
As already anticipated in the introduction, we bring two conceptual improvements to the original recipe of Ref. 3. First, we define a physics-informed collective variable (CV) $z(\mathbf{x})$ that is based on the committor and is effective when used with metadynamics-like enhanced sampling approaches. Second, we improve such approaches by combining them with the Kolmogorov bias potential to favor the TS sampling.
@@
Even if the committor has often been thought to be an ideal CV, Berezhkovskii and Szabo (2005); Li and Ma (2014); Ma and Dinner (2005); He, Chipot, and Roux (2022) we already pointed out in the introduction that its straightforward application in an enhanced sampling context is far from optimal due to the high degeneracy of the metastable states along such a variable, as $q(\mathbf{x_{A}})\simeq 0$ and $q(\mathbf{x_{B}})\simeq 1$, and to its sharp behavior in the transition region. This would suggest that a smoother modification of the committor might be an effective CV. A practical way of simply and economically achieving this result comes at zero cost from the NN-based parametrization of the committor proposed in Ref. 3. There, as we have seen in Section IV.2 and depicted in Figure 1, the committor is represented as a NN, and a sharp sigmoid activation function $\sigma$ is applied to the output layer of such a network to enforce the correct functional form to the learned function, that is, $q(\mathbf{x})=\sigma(z(\mathbf{x}))$. It is thus wiser to employ as a CV, not the activated final committor function, but rather the output $z(\mathbf{x})$ of the NN. In fact, this still preserves the physical information encoded in the committor function, but it is also suitable to be used in an enhanced sampling context as it still has finite gradients in the metastable regions, and it is not as sharp in the TS region.
@@
As one can expect, similar numerical issues could also affect the bias potential $V_{K}$ that, in Ref. 3, was computed as a function of the gradients of the committor function $\nabla q(\mathbf{x})$. In fact, this choice is not optimal in practice as, due to the sharp behavior of $q(\mathbf{x})$, such a computation can easily lead to extremely small values below machine precision in the basins where $q(\mathbf{x})$ is flat (see Supplementary Figure S1), thus making the bias almost ineffective there. Conveniently, by taking advantage of the change of variable $q\rightarrow z$ and with the help of some elementary algebra (see Supplementary Section S1), one can bypass this issue, writing the bias potential as a function of the smoother and more numerically stable gradients of $z$.
@@
Like any CV, $z$ can be used, in principle, with any CV-based method. However, here, we opt for the On-the-fly Probability Enhanced Sampling (OPES) method Invernizzi and Parrinello (2020); Trizio et al. (2024), a recent evolution of metadynamics Laio and Parrinello (2002) that arguably represents the state of the art in this field. We already pointed out in Ref. 3 that, in metadynamics-like frameworks, the metastable basins are filled with an external bias potential to reduce the height of free energy barriers and promote transitions, but the TS-region is still found at a local maximum, whose sampling is thus unfavored (see Figure 1).
@@
As one can expect, the improved sampling of this new approach is also beneficial to the iterative procedure as it greatly simplifies the dataset construction, exploiting the best from the two biasing schemes. From the Kolmogorov bias, it inherits an extensive sampling of the TSE, which makes it possible to apply accurately the variational principle (see Section IV.1). On the other hand, from OPES, we obtain a fast-converging bias potential that allows sampling of the relevant part of the phase space. Invernizzi and Parrinello (2022); Trizio et al. (2024) This facilitates the reweighting, thus guaranteeing a more accurate estimate of the Boltzmann weights and substantially reduces the number of iterations needed for convergence (see Section IV.2). Moreover, integrating eventual additional CVs in the OPES biasing scheme is simple, which could help to further improve the simulation efficiency in some complex cases, for instance, biasing orthogonal degrees of freedom that distinguish different paths if present.
IV.4 Summary of self-consistent iterative procedure
@@
  • Step 1: The NN-based committor $q_{\theta}^{n}(\mathbf{x})$ at iteration $n$ is optimized on the dataset of configurations $\mathbf{x}^{n}$ and weights $w_{i}^{n}$ updated from the previous iterations. To start the procedure, at the first iteration $n=0$, we build such a dataset by running some (short) simulations in the metastable basins and labeling them accordingly. The optimization is based on the variational principle of Eq. 5, minimizing the loss function of Eqs. 8 and 9 (see Section IV.2).
  • Step 2: We perform biased simulations under the coupled action of the Kolmogorov and OPES bias potentials, $V^{n}_{K}$ and $V^{n}_{\text{OPES}}$. The latter is applied along the committor-related CV $z^{n}(\mathbf{x})$ and drives the sampling across the whole relevant configurational space between the two states, whereas the former guarantees a thorough sampling of the TS region crucial for applying Eq. 5 (see Section IV.3).
@@
Data Availability
Training and simulation data and inputs are available on GitHub tri and on ZenodoTrizio, Kang, and Parrinello (2025)
Code Availability
The code for the training of the NN-based committor model alongside didactic tutorials is available through the open-source mlcolvar library Bonati et al. (2023), which is the preferred way to access the most updated code. To obtain the results reported in the manuscript, version 1.2.2 was used and a frozen version is also available on ZenodoTrizio, Kang, and Parrinello (2025) The PLUMED Tribello et al. (2014); consortium (2019) interface for the application of the bias is available on GitHub tri and on ZenodoTrizio, Kang, and Parrinello (2025)
Acknowledgements.
The Authors are grateful to Luigi Bonati, Andrea Rizzi, Jintu Zhang, Umberto Raucci, Alice Triveri, and Francesco Mambretti for the useful discussions and feedback about this manuscript.